A monopolist faces two totally separated markets with inverse demand p=100 – qA and
p=160−2qB respectively. The monopolist has no fixed costs and a marginal cost given by mc= 2 /3
q Find the profit maximizing total output and how much of it that is sold on market A and market
B respectively if the monopoly uses third degree price discrimination.
a) What prices will our monopolist charge in the two separate markets? (6 m)
b) Calculate the price elasticity of demand in each market and explain the intuition behind the
relationship between the prices and elasticities in these two separate markets
1
Expert's answer
2020-06-12T12:03:05-0400
"MR=TR^\/"


"TR=\displaystyle\sum_{i=1}^np_iq_i"


"TR=100q-q^2+160q-2q^2"


"TR=260q-3q^2"


"MR=260-6q"


"MR=MC"


"260-6q=\frac{2}{3}q"


"q=39"

a) If the market is a monopoly, then "mc=p_i"


For A:


"\frac{2}{3}q=100-2q"


"\frac {8}{3}q=100"

"q=36.5"


"p=63.5"

For B:


"\frac{2}{3}q=160-4q"

"\frac {14}{3}q=160"


"q=34.2"


"p=91.6"

b)

"E_A=\frac{\varDelta Q_A}{\varDelta p_A}\times \frac{p}{q}=-0.575"


"E_B=\frac{\varDelta Q_B}{\varDelta p_B}\times \frac{p}{q}=-\frac{1}{2} \times \frac{2}{3}=-0.325"

The sum of the modules of the elasticity coefficients of each product individually is 1